OpenStudy (anonymous):

Solve 7x - 9 - 12x = 26. -7 -2 2 7

OpenStudy (jiteshmeghwal9):

7x-9-12x=26 -5x-9=26 -5x=26+9 -5x=35 x=35/-5=-7.

OpenStudy (anonymous):

7x - 9 - 12x = 26. subtract 7x-12x to get -5x -5x-9=26 add 9 to each side -5x=35 divide -5 from each side x= -7